OpenStudy (anonymous):

What is 43/9 as a mixed number

OpenStudy (aaliyahfairgood):

4+(7/9))(3/5) = ((36/9)+(7/9))(3/5) = ((36+7)/9)(3/5) = (43/9)(3/5) = (43*3)/(9*5) = 43/15 = (30+13)/15 = (30/15) + (13/15) = 2+(13/15) 0 (4+(7/9))(3/5) = ((36/9)+(7/9))(3/5) = ((36+7)/9)(3/5) = (43/9)(3/5) = (43*3)/(9*5) = 43/15 = (30+13)/15 = (30/15) + (13/15) = 2+(13/15)

OpenStudy (danjs):

How many times can 9 go in to 43 total?

OpenStudy (danjs):

that is your whole number, the remaining amount is a fraction of something over 9

OpenStudy (danjs):

\[4 + \frac{ 7 }{ 9 }\]

OpenStudy (aaliyahfairgood):

it goes in 4 times

OpenStudy (danjs):

and you are left with 7 remaining

OpenStudy (aaliyahfairgood):

with a remainder of 7 4(whole number) 7/9

OpenStudy (danjs):

yes

OpenStudy (danjs):

you can write it like this \[\frac{ 43 }{ 9 } = \frac{ 36 + 7 }{ 9 } = \frac{ 36 }{ 9 } + \frac{ 7 }{ 9 }\]

OpenStudy (danjs):

4 and 7/9ths

OpenStudy (aaliyahfairgood):

that is better

OpenStudy (anonymous):

Close this question, you have it answered..